10-5a CAPM Appronch Calculate the cost of equity financing given the following Risk-free ratc 1% Market risk premium: 7% 1.25 Beta: 10-5b Bond-Yield-Plus-Risk-Premium Approsch The return on a bond is 4% while the return on common is 5%. What is the risk premium? 10-5e DCE Approach Solve for required return given the following The dividend paid at the beginning of the first time period (Du) was fa/share. The dividend grows in perpetuity at 3.5% (growth rate g). The price of the stock (M) is $39/share. 10-5d Averaging the Alternative Estimates The average rate of return on a stock was calculated to be 18.4%. The DCF method yielded 21% while the te-RP method yielded i 4%. What estimate did the CAPM yield? 10-6 Cost of New Common Stock Can a firm expect capital from issuance of new shares to cost more or less than using retained eamings? Explain your answer 10-7 Composite WACC Calculate the WACC (report using XX% format) given the follow information, assuming a tax rate of 30%; 20% 7% woc 20% 8% 60% 12%
Answer #1
10-5a) Cost of equity per CAPM = risk free rare+beta*market risk premium = 1%+1.25*7% = 9.75%
10-5b) Risk premium = return on common - return on bond = 5% - 4% = 1.00%
10-5c) Required return per DCF (constant dividend growth) = D0*(1+g)/P0+g = 4*1.035/39+0.035 = 14.12%
10-5d) We have 18.4% = (21%+14%+rcapm)/3
rcapm = 18.4%*3-21%-14% = 20.20%
10-6) The cost of new shares will be more than the cost of retained earnings as new equity has flotation cost.
10-7) After tax cost of debt = 7%*(1-30%) = 4.90%
WACC:
Component Cost Weight WACC
4.90% 20% 0.98%
8.00% 20% 1.60%
12.00% 60% 7.20%
WACC 9.78%